calculate the molarity of a 45.3g NaNo³ dissolved in water to produce 225 ml of solution


1
Expert's answer
2021-04-13T08:04:16-0400

Molar Mass of NaNO3 = 84.9947

= 45.3/84.9947 = 0.533 moles

= 45.3+ 225 = 270.3

= 0.2703 L

= 0.533/0.2703

= 1.972mol/L